OpenStudy (jinx44):

Which ordered pairs are solutions to the equation? y = 3x + 1 A. (6, 19) and (0, 1) B. (19, 6) and (0, 1) C. (6, 18) and (0, 1) D. (6, 19) and (0, 0) will fan and medal if answer is right!

OpenStudy (jinx44):

plz i am timed =(

rebeccaxhawaii (rebeccaxhawaii):

plug it in

OpenStudy (anonymous):

ill help u do k12?

OpenStudy (jinx44):

@rebeccaxhawaii ?

OpenStudy (jinx44):

oh ok @CHESSJOB

OpenStudy (anonymous):

its D since you plugin the coordinates into the equation

OpenStudy (jinx44):

ok could u explain how u 'plug them in'?

OpenStudy (anonymous):

shoot

OpenStudy (anonymous):

wait

OpenStudy (jinx44):

mk

OpenStudy (anonymous):

its actually A

OpenStudy (jinx44):

yea @KamiBug told me that in message.

OpenStudy (jinx44):

lol

OpenStudy (anonymous):

y=3x+1 A) 19=3(6) +1 19 = 19 B) 1= 0 + 1 1 = 1

OpenStudy (jinx44):

thanks again @KamiBug and @CHESSJOB!